Operation Summary for August 17, 2011 - August 19, 2011 |
Wednesday August 17 The day shift began with Operations monitoring the Tevatron (TEV) store 8985 (92.45E30) and the Recycler stash (393.3E10), with the Antiproton Source (Pbar) stacking (28.1E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, EE Support personnel began work at 8:19 AM on two Switchyard (SY) power supplies (S:QP3 and S:HP3DS). Operations terminated store 8985 at 10:13 AM. Operations put the TEV into shot setup for a 3x3 study store and established store 8986 at 11:22 AM with an average initial luminosity (AIL) of 25.83E30; this study is a destructive beam study and not useful to the experiments. FESS personnel cleaned the strainer for the TEV sector F3 heat exchanger, at 1:54 PM, which improved the flow of pond water. A Booster RF station (BRF16) began to repeatedly trip off; RF personnel investigated but could not find the problem. They thought the problem was in the tunnel; they will bring BRF1 online. At 3:02 PM, the TEV quenched at D0 due to bad beam positions. Operators prepared D0 and CDF for access at 3:26 PM. Booster came back online at 3:35 PM. During the evening shift, D0 and CDF came out of access at 4:10 PM. Operators put the TEV through a dry squeeze at 5:32 PM and through at wet squeeze 6:13 PM. At 6:50 PM, the Pelletron suffered a full discharge; operators put it into recovery mode. Operations put the TEV into shot setup and established store 8990 at 8:15 PM with an average initial luminosity (AIL) of 334.78E30. Thursday August 18 The midnight shift began with Operations monitoring store 8990 (167.36E30) and the stash (140.5E10), with Pbar stacking (132.5E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 1:28 AM, the MI ramp tripped off; it reset. At 1:29 AM, a TEV RF station (TRF7) tripped off; operators will turn it back on in between stores. At 1:46 AM, the beam permit tripped off for the P1 and P2 transfer lines and a Pbar Debuncher power supply (D:HT727) tripped off; operators replaced a fuse for the power supply. Operations terminated store 8990 at 2:39 AM and then prepared CDF for access. BRF5 tripped off for the first of two times at 2:55 AM; after the second trip at 4:10 AM, it would not reset; operators bypassed the station. CDF came out of access at 6:33 AM. Operators put the TEV through dry and wet squeezes. During the day shift, Operations put the TEV into shot setup at 8:40 AM. At 8:47 AM, operators reported seeing a power glitch that caused a Linac Klystron RF station (KRF7), MI kickers, and Pbar sweeper magnets to trip off. The High Voltage Coordinator reported a ComEd glitch off site. Operators dealt with many other problems due to this power glitch. Operators put the TEV through a dry squeeze at 10:31 AM and through a wet squeeze at 11:25 AM. At 12:34 PM, operators reported losing a transfer of antiprotons from Pbar to the Recycler due to a MI beam permit trip. Operations put the TEV into shot setup at 1:07 PM. At 1:25 PM, operators reported problems loading protons due to Booster beam permit trips and MI damper problems; the MI Operations Specialist contacted a damper person and then made adjustments. Operations established store 8997 at 2:46 PM with an AIL of 333.71E30. Operators prepared enclosure MI-31 and the Pelletron for access at 3 PM. Operations reported receiving a signed Operational Readiness Clearance for a new Meson FTBF experiment T-1015 at 3:52 PM. During the evening shift, at 4:19 PM, a NuMI kicker (I:KPS6NF) power supply tripped off and took down the NuMI beam permit; everything reset, but then the NuMI, Pbar, and MI beam permits began to trip on a MECAR indication; Operations set MECAR to relearn and the trips stopped. Operations established beam to T-1015 at 5:56 PM and disabled their beam at 8:51 PM. Operators reported a vacuum burst at MI-60 at 11:33 PM; it recovered. Friday August 19 The midnight shift began with Operations monitoring store 8997 (96.47E30) and the stash (157.5E10), with Pbar stacking (34.0E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8997 at 3:17 AM and put the TEV into shot setup. At 3:27 AM, LRF2 tripped off for the first of three times in 16 minutes; it reset. At 3:57 AM, LRF1 tripped off; it reset. At 4:07 AM, LRF1 and LRF2 tripped off; LRF1 reset, but operators investigated and had to lower the voltage for the LRF2 power amplifier to get it operating normally. Operators reported that shot setup had to be extended due to continuing problems with LRF1 & 2. Operations established store 8998 at 5:14 AM with an AIL of 121.45E30; the low luminosity was due to the Pelletron being off. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow TEV personnel to conduct a 3X3 study store during the day today and will also allow MI and Booster personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 08/19/2011 email Fermilab |
| Security, Privacy, Legal | |